the latus rectum subtends a right angle at the centre of the ellipse then its eccentricity is

Grade:11

2 Answers

Vijay Mukati
8 years ago
Dear Student,

Coordinates of end points of LR will be (ae , +/- b2/a). Since it substends right angle at cenre therefore, the produc of slopes of this two points from certre = -1
m1 * m2 = -1 or
(b2/a2e) * (-b2/a2e) = -1
Therfore, e = b2/a2

Thanks
NARAYANA
8 years ago
IN CINIC SECTIONS THESE TYPE OF QUESTION ARE LOT. TO DO THESE PROBLEMS LOT OF PRACTISE IS REQUIRED.SO ANSWER OF YOUR QUESTION IS c=b2/a2.